OpenStudy (anonymous):

The area of a rectangle is x^2 - 2x - 15 and the length of the rectangle is x + 3. Find the width of the rectangle. A. x – 5 B. x2 + 15 C. x + 5 D. (x – 2)(x + 3)

OpenStudy (anonymous):

http://www.youtube.com/watch?v=Ldsc_bJG1Ic

OpenStudy (anonymous):

the area=length* width x^2-2x-15=(x+3)(width) (x-5)(x+3)=(x+3)(width) Width=(x-5)(x+3)/(x+3) Width=x-5 the answer is A

OpenStudy (anonymous):

@ktnguyen1 thanks